OpenStudy (anonymous):

State whether the slope of a line passing through (2, -3) and (1, 8) is positive, negative, zero or undefined

OpenStudy (anonymous):

It isn't zero, that would mean the y values would be the same. It isn't undefined, that would me the x values would be the same. It must be positive or negative. The slope is rise/run, or (change in y)/(change in x). In this case, it is: (8- (-12))/ (8-6) = 20/2 = 10, since this is a positive number, the slope is positive.